OpenStudy (anonymous):

x-y+2z=6 2x+y+z=3 x+y+z=2

OpenStudy (anonymous):

add 1 and 2 3x + 3z = 9 x + z = 3..................a add 1 and 3 2x + 3z = 8...............b solvin a and b multiply a by 2: 2x + 2z = 6...............c b - c:- z = 2 plug z=2 in b: 2x + 6 = 8 x = 1 plug x=1 and z = 2 in first equation 1-y + 4 = 6 y = 1+ 4 - 6 = -1 solution is x=1, y = -1 and z = 2

OpenStudy (anonymous):

ok with that ? the results check out ok